# Rethinking Prompt Optimizers: From Prompt Merits to Optimization

### v2

@@ Abstract
Prompt optimization (PO) provides a practical way to improve response quality when users lack the time or expertise to manually craft effective prompts. Existing methods typically rely on advanced, large-scale LLMs like GPT-4 to generate optimized prompts. However, due to limited downward compatibility, verbose, and instruction-heavy prompts from advanced LLMs can overwhelm lightweight inference models and degrade response quality. In this work, we rethink prompt optimization through the lens of explicit and interpretable design. We first identify a set of model-agnostic prompt quality merits and empirically validate their effectiveness in enhancing prompt and response quality. We then introduce a merit-guided prompt optimizer (MePO), which is locally deployable and trained on our preference dataset built from merit-guided prompts generated by a lightweight LLM. Unlike prior work, MePO avoids online optimization reliance, reduces cost and privacy concerns, and—by learning clear, interpretable merits—generalizes effectively to both large-scale and lightweight inference models. Experiments demonstrate that MePO achieves better results across diverse tasks and model types, offering a scalable and robust solution for real-world deployment. The code and dataset can be found in https://github.com/MidiyaZhu/MePO.
@@ 1 Introduction
One line of work uses API-based models such as GPT-4 for online prompt optimization, typically inserting a task-specific meta-prompt before or after each query to reduce cost, while leaving the query itself unoptimized (Xiang et al., 2025; Guo et al., 2024; Ye et al., 2024). However, this approach requires prior task knowledge, limiting its applicability in open-ended, real-world scenarios. Another line explores locally deployable optimizers that optimize each prompt and are trained independently of specific tasks, though their training data still originates from prompts generated by advanced online LLMs, implicitly treating these outputs as optimal (Lu et al., 2025; Cheng et al., 2024).
Despite the differences in deployment strategies, both online and local optimization approaches rely on prompts produced by advanced online LLMs as the gold-standard optimal prompts. This design implicitly assumes downward compatibility: the notion that prompts generated by large-scale LLMs can be effectively interpreted by lightweight inference models. However, our empirical analysis challenges this assumption. As shown in Fig. 1(a), prompts crafted by advanced, large-scale LLMs often contain verbose, instruction-heavy content, which may overwhelm lightweight models, causing them to miss the key point or engage in excessive reasoning, ultimately leading to incorrect or irrelevant responses.
@@
Lightweight models are typically not used for prompt optimization due to their perceived limitations in handling complex optimization processes Wang et al. (2024); Lu et al. (2025), as further illustrated in Fig. 1(b). However, our research reveals a promising alternative. We find that these models can indeed serve as effective prompt optimizers when guided by explicit and interpretable prompt design merits (Fig. 1(c)). This suggests that, with clear, learnable structures, even lightweight LLMs can function as capable optimizers, offering a cost-efficient and locally deployable alternative to online methods, particularly in low-resource or privacy-sensitive settings.
To this end, we propose MePO, a Merit-Guided Prompt Optimization model. We begin by conducting a systematic empirical analysis to identify four sets of interpretable prompt merits that consistently characterize high-quality prompts, offering actionable insights into effective prompt construction. Guided by these merits, we build a prompt optimization preference dataset using optimized prompts generated by a lightweight LLM—eschewing reliance on advanced online LLMs as in prior work—and use it to train an end-to-end prompt optimization model. At the prompt level, our approach produces precise and clear merit guidance that can be generated by lightweight prompt optimization models. At the response level, our optimized prompts not only enhance the performance of similarly scaled inference models, but also demonstrate strong downward and upward compatibility, as the learned prompt merits generalize effectively across both large-scale and lightweight inference models. Besides, constructing the dataset using lightweight models greatly reduces API cost and external dependencies, enabling scalable, privacy-preserving, and locally deployable prompt optimization.
@@ 2 Empirical Analysis: What Merits a Good Prompt?
We argue that these limitations are not solely due to the generation capacity of lightweight LLMs, but rather to the overly vague or underspecified optimization instructions used in prior work. Specifically, Wang et al. (2024) adopt the prompt “Given error feedback, generate a better prompt,” and Lu et al. (2025) use “You are an expert in prompt optimization.” Both are overly general and lack concrete guidance. In contrast, we find that when lightweight LLMs are explicitly instructed with detailed prompt patterns, they can generate effective optimized prompts despite their limited capacity (Fig. 1(c)). This observation motivates a key question: What merits a good prompt?
@@ 2.1 Merit Discovery
As shown in Fig. 3, prompts optimized using our designed merits—despite being applied to a lightweight model—can match or even outperform GPT-3.5-generated responses (Alpaca) and significantly beat human-preferred responses (BPO).
Prompt-Level Evaluation: To move beyond response quality and directly assess prompt quality—a dimension largely overlooked in prior work—we randomly interleaved optimized and raw prompts and asked DeepSeek-R1 to evaluate them in isolation (i.e., without seeing the generated responses). Across 2000 prompt comparisons per dataset, optimized prompts were preferred in 95.75% of cases on BPO and 97.9% on Alpaca. These results reinforce the effectiveness of our proposed prompt merits and demonstrate that prompt quality can be reliably evaluated independently of generated outputs.
Summary: Our evaluations demonstrate that optimized prompts grounded in our proposed interpretable merits not only yield stronger responses but also exhibit higher prompt quality. Moreover, the results show that effective prompt optimization is not exclusive to advanced, large-scale LLMs such as GPT-4. With merit-based guidance, lightweight LLMs can also serve as capable prompt optimizers, offering scalable and privacy-friendly alternatives for real-world deployment.
The detailed implementation of merit discovery process is provided in Appx.B.
@@ 3.1 Constructing the Merit-Aligned Prompt Preference Dataset
Since Alpaca and BPO contain relatively well-formed prompts—either produced by text-davinci-003 or curated by humans—we further simulate real-world usage by randomly selecting 10% of $P_{\text{silver}}$ and intentionally degrading them using a base model. These lower-quality prompts better reflect common user inputs and are paired with their corresponding $P_{\text{golden}}$ to enhance data diversity.
@@
# 4-Tuples 25,526 3,000 10,225 1,400
Total samples: 40,151
Validation (Win Rate) GPT-4o (5k) DeepSeek-R1 (40k)
Response Win Rate 94.82% 100.00%
Prompt Win Rate 97.68% 97.62%
Table 1: Overview of our POP dataset. The top section reports the number of prompt–response 4-tuples derived from Alpaca and BPO. The bottom section shows validation results from DeepSeek-R1 and GPT-4o, with win rate denoting the proportion where $R_{\text{golden}}$ or $P_{\text{golden}}$ outperforms its silver counterpart.
@@
Qwen2-7b - 80.68 65.19 76.88 45.42 80.20 69.67
Inference Model 81.22 66.21 79.01 48.66 81.34 71.29
BPO 80.89 66.38 77.38 50.07 82.64 71.47
FIPO 82.37 67.49 82.71 49.38 82.48 72.89
MePO 83.33 68.52 83.12 53.42 83.46 74.37
Tulu2-7b-dpo - 45.37 26.54 30.63 32.68 70.02 41.05
Inference Model 46.40 29.35 32.93 35.13 70.78 42.92
BPO 47.98 31.57 32.08 35.96 71.60 43.84
FIPO 50.55 36.95 33.66 36.06 72.63 45.97
MePO 55.05 38.14 35.18 39.89 73.60 48.37
LLaMA2-7b-chat-hf - 35.40 29.27 15.06 31.08 49.95 32.15
Inference Model 36.74 29.52 16.27 33.49 51.41 33.47
BPO 38.30 30.89 19.85 36.34 52.56 35.59
FIPO 36.24 29.44 17.03 35.81 51.58 34.02
MePO 39.86 31.74 25.25 38.68 55.33 38.17
LLaMA3-8b-instruct - 46.72 41.31 55.04 35.21 65.56 48.77
Inference Model 52.02 42.32 57.87 37.83 67.19 51.45
BPO 50.88 44.37 58.39 39.33 71.81 52.96
FIPO 49.12 44.97 56.81 39.29 75.56 53.15
MePO 55.01 51.88 59.59 42.19 76.66 57.07
Table 2: Performance comparison of prompt optimization methods across multiple datasets and inference models. ‘-’ denotes the original unoptimized prompt. ‘Inference Model’ refers to using the inference model itself as the optimizer (based on the instruction format in Fig. 11). Detailed BBH task-level results are provided in Appx.D.3.
After filtering and aggregation, the final POP dataset comprises 40k samples. Table 1 summarizes the dataset composition and presents quality validation scores from both DeepSeek-R1 and GPT-4o, showing consistently strong preference for the optimized prompts across both response- and prompt-level evaluations.
@@ 3.2 Training the Merit-Guided Prompt Optimizer
[…]
@@ 4 Experiments
We evaluate the effectiveness of MePO from the following perspectives: (1) Comparison with SOTA local prompt optimizers (Sec. 4.2); (2) Ablation study with training-free prompting (Sec. 4.3); (3) Upward compatibility in larger inference models (Sec. 4.4); (4) Comparison with online-based meta-prompts (Sec. 4.5) and (5) Effectiveness analysis of optimized prompts (Sec. 4.6).
@@ 4.1 Setups
MePO is trained using Qwen2.5-7B-Instruct2, selected for its strong generation ability among similarly sized models Qwen et al. (2025). This choice ensures that the optimizer is built on a lightweight LLM that maximizes performance within its size class—allowing efficient deployment without sacrificing generation quality.
Datasets We evaluate our method on five benchmark datasets: ARC-Easy, ARC-Challenge (Clark et al., 2018), GSM8K (Cobbe et al., 2021), BigBench-Hard (BBH) (Suzgun et al., 2023), and PiQA (Bisk et al., 2020). Following Lu et al. (2025), we adopt 3-shot for GSM8K, BBH, and PiQA, and zero-shot for ARC, with instructions for answer extraction. Prompt optimization is applied only to test query, while others remain unchanged.
Baselines We compare MePO with two state-of-the-art local prompt optimization models: BPO (Cheng et al., 2024) and FIPO (Lu et al., 2025). We evaluate optimizers’ performance across three inference model families: Qwen (Yang et al., 2024), Tulu (Ivison et al., 2023) and LLaMA (Touvron et al., 2023). We also use the inference model itself as an optimization baseline, motivated by the inherent alignment between a model and its own preferences, as leveraging an LLM’s natural capability to refine prompts for itself better reflects its aligned outputs (Xiang et al., 2025).
Additional implementation details and dataset descriptions are provided in Appx.D.
@@ 4.2 Main Results
We present the performance of each optimizer, including our proposed MePO, on different inference models across five datasets. The results, summarized in Table 2, lead to the following conclusions:
Efficacy MePO consistently outperforms all baselines across datasets and task models, with average accuracy gains of 1% on Qwen2-7B, 2.37% on Tulu2-7B, 2.59% on LLaMA2-7B, and 3.91% on LLaMA3-8B. Notable improvements are observed on ARC-Challenge with LLaMA3-8B (+6.91%), GSM8K with LLaMA2-7B (+5.4%), and ARC-Easy and BBH with Tulu2-7B (+4.5%, +3.69%).
Robustness Inference models often perform better when the optimizer shares the same architecture, reflecting alignment with model-specific preferences. For example, in the BPO–FIPO comparison, BPO (based on LLaMA2) performs better when responding with LLaMA2-7B, while FIPO (based on Tulu2) excels with Tulu2-7B. However, both optimizers show misalignment on models like LLaMA3 and Qwen2. On ARC-Easy, inference model optimization outperforms BPO (+1.14% with LLaMA3, +0.33% with Qwen2) and FIPO (+2.90% with LLaMA3). Similarly, on GSM8K, the inference model surpasses BPO (+1.63% with Qwen2) and FIPO (+1.06% with LLaMA3). In contrast, MePO consistently outperforms inference model optimization across all datasets, indicating that its learned prompt merits are broadly interpretable across architectures, demonstrating both robustness and strong generalization capability.
Task Model Optimizer ARC-Easy ARC-Challenge GSM8K BBH PiQA Avg.

### v3

@@ Abstract
Prompt optimization (PO) provides a practical way to improve response quality when users lack the time or expertise to manually craft effective prompts. Existing methods typically rely on LLMs’ self-generation ability to optimize prompts. However, due to limited downward compatibility, the instruction-heavy prompts generated by advanced LLMs can overwhelm lightweight inference models and degrade response quality, while also lacking interpretability due to implicit optimization. In this work, we rethink prompt optimization through the lens of explicit and interpretable design. We first identify a set of model-agnostic prompt quality merits and empirically validate their effectiveness in enhancing prompt and response quality. We then introduce MePO, a merit-guided, locally deployable prompt optimizer trained on our merit-guided prompt preference dataset generated by a lightweight LLM. MePO avoids online optimization, reduces privacy concerns, and, by learning clear, interpretable merits, generalizes effectively to both large-scale and lightweight inference models. Experiments demonstrate that MePO achieves better results across diverse tasks and model types, offering a scalable and robust solution for real-world deployment. The code, model and dataset can be found in https://github.com/MidiyaZhu/MePO.
@@ 1 Introduction
One line of work uses API-based models such as GPT-4 for online prompt optimization, typically inserting a task-specific meta-prompt before or after each query to reduce cost, while leaving the query itself unoptimized (Xiang et al., 2025; Ye et al., 2024). This approach requires prior task knowledge, limiting its use in open-ended scenarios. Another line explores locally deployable optimizers that optimize each prompt and are trained independently of specific tasks. However, most of their training data still originates from prompts generated by advanced online LLMs (Lu et al., 2025; Cheng et al., 2024), or selecting optimized prompts based on the LLMs’ implicit self-generation and evaluation capabilities (Liang et al., 2024).
Despite the differences in deployment strategies, both online and local optimization approaches rely on prompts produced by advanced online LLMs as the gold-standard optimal prompts. This design implicitly assumes downward compatibility: the notion that prompts generated by large-scale LLMs can be effectively interpreted by lightweight inference models. However, our empirical analysis challenges this assumption. As shown in Fig. 1(a), prompts from large-scale LLMs are often verbose and instruction-heavy, which may overwhelm lightweight models with limited memory and reasoning capacity Yang et al. (2023), leading to missed key points or excessive reasoning and ultimately incorrect or irrelevant responses.
@@
Lightweight models are typically not used for prompt optimization due to their perceived limitations in handling complex optimization processes Wang et al. (2024); Lu et al. (2025), as further illustrated in Fig. 1(b). However, our research reveals a promising alternative. We find that these models can indeed serve as effective prompt optimizers when guided by explicit and interpretable prompt design merits (Fig. 1(c)). This suggests that, with clear, learnable structures, even lightweight LLMs can function as capable optimizers, offering an efficient and locally deployable alternative to online methods, particularly in low-resource or privacy-sensitive settings.
To this end, we propose MePO, a Merit-Guided Prompt Optimization model. Unlike implicit optimization like POIR Liang et al. (2024), we begin by conducting a systematic analysis to identify four sets of explicit prompt merits that consistently characterize high-quality prompts, offering actionable insights into effective prompt construction. Guided by these merits, we build a prompt optimization preference dataset using optimized prompts generated by a lightweight LLM, eschewing reliance on advanced online LLMs, and use it to train an end-to-end prompt optimization model. At the prompt level, our approach produces precise and clear merit guidance that can be generated by lightweight prompt optimization models. At the response level, our optimized prompts not only enhance the performance of similarly scaled inference models, but also demonstrate strong downward and upward compatibility, as the learned prompt merits generalize effectively across both large-scale and lightweight inference models. Besides, using lightweight models for dataset construction reduces external dependencies, enabling scalable, privacy-preserving local deployment.
@@ 2 Empirical Analysis: What Merits a Good Prompt?
We argue that these limitations are not solely due to the generation capacity of lightweight LLMs, but rather to the overly vague or underspecified optimization instructions used in prior work. Specifically, Wang et al. (2024) adopt the prompt “Given error feedback, generate a better prompt”, and Lu et al. (2025) use “You are an expert in prompt optimization”. Both are overly general and lack concrete guidance. In contrast, we find that when lightweight LLMs are explicitly instructed with detailed prompt patterns, they can generate effective optimized prompts despite their limited capacity (Fig. 1(c)), which aligns with findings that lightweight LLMs—with limited contextual memory Yang et al. (2023)—particularly benefit from explicit and structurally clear prompts Ying et al. (2024). This observation motivates a key question: What merits a good prompt?
@@ 2.1 Merit Discovery
As shown in Fig. 3, prompts optimized using our designed merits, despite being applied to a lightweight model, can match or even outperform GPT-3.5-generated responses (Alpaca) and significantly beat human-preferred responses (BPO).
Prompt-Level Evaluation: To move beyond response quality and directly assess prompt quality, a dimension largely overlooked in prior work, we asked DeepSeek-R1 to evaluate raw and optimized prompts in isolation. Across 2000 prompt comparisons per dataset, optimized prompts were preferred in 95.75% of cases on BPO and 97.9% on Alpaca. These results reinforce the effectiveness of our proposed prompt merits and demonstrate that prompt quality can be reliably evaluated independently of generated outputs.
Summary: Our evaluations demonstrate that prompts optimized using these interpretable merits yield higher prompt quality and stronger responses. Moreover, effective prompt optimization is not exclusive to advanced, large-scale LLMs; with merit-based guidance, lightweight LLMs can also serve as capable and practical prompt optimizers.
The detailed implementation of the merit discovery process is provided in Appx.B.
@@ 3.1 Constructing the Merit-Aligned Prompt Preference Dataset
Since Alpaca and BPO contain relatively well-formed prompts, either produced by text-davinci-003 or curated by humans, we further simulate real-world usage by randomly selecting 10% of $P_{\text{silver}}$ and intentionally degrading them using a base model. These lower-quality prompts better reflect common user inputs and are paired with their corresponding $P_{\text{golden}}$ to enhance data diversity.
@@
# 4-Tuples 25526 3000 10225 1400
Total samples: 40151
Validation (Win Rate) GPT-4o(5k) DeepSeek-R1(40k) Human (400)
Response Win Rate 94.82% 100.00% 81.50%
Prompt Win Rate 97.68% 97.62% 83.99%
Table 1: Overview of our POP dataset. The top section reports the number of prompt–response 4-tuples derived from Alpaca and BPO. The bottom section shows validation results from DeepSeek-R1, GPT-4o and four human evaluations, with win rate denoting the proportion where $R_{\text{golden}}$ or $P_{\text{golden}}$ outperforms its silver counterpart.
@@
Qwen2-7b - 80.68 65.19 76.88 49.34 80.20 70.46
Inference Model 81.22 66.21 79.01 52.68 81.34 72.09
BPO 80.89 66.38 77.38 53.07 82.64 72.07
FIPO 82.37 67.49 82.71 52.74 82.48 73.56
MePO 83.33 68.52 83.12 54.35 83.46 74.56
Tulu2-7b-dpo - 45.37 26.54 30.63 35.67 70.02 41.65
Inference Model 46.40 29.35 32.93 38.27 70.78 43.55
BPO 47.98 31.57 32.08 39.33 71.60 44.51
FIPO 50.55 36.95 33.66 39.56 72.63 46.67
MePO 55.05 38.14 35.18 43.25 73.60 49.04
LLaMA2-7b-chat-hf - 35.40 29.27 17.51 34.17 49.95 33.26
Inference Model 36.74 29.52 18.88 36.65 51.41 34.64
BPO 38.30 30.89 25.25 39.60 52.56 37.32
FIPO 36.24 29.44 24.72 39.14 51.58 36.22
MePO 39.86 31.74 29.42 41.97 55.33 39.66
Gemma2-9b - 89.86 38.05 63.03 61.41 82.01 66.87
Inference Model 90.61 40.36 64.75 64.29 82.54 68.51
BPO 91.41 38.16 65.36 64.30 83.04 68.45
FIPO 89.88 46.29 64.75 66.10 83.35 70.07
MePO 92.30 48.89 68.67 69.47 85.35 72.94
Qwen3-8b - 82.11 52.60 89.46 68.07 77.80 74.01
Inference Model 83.85 52.72 90.07 69.45 79.87 75.19
BPO 82.75 53.12 86.96 70.03 79.11 74.39
FIPO 78.53 50.11 83.55 70.21 73.72 71.22
MePO 84.38 54.68 91.12 71.12 83.08 76.88
Table 2: Performance comparison of prompt optimization methods across multiple datasets and inference models. ‘-’ denotes the original unoptimized prompt. ‘Inference Model’ refers to using the inference model itself as the optimizer (based on the instruction format in Fig. 11). Detailed BBH task-level results are provided in Appx.D.1.
After filtering and aggregation, the final POP dataset comprises 40k samples. Table 1 summarizes the dataset composition and quality validation scores from both advanced LLMs and humans, showing consistently strong preference for the optimized prompts across both response- and prompt-level evaluations. More analysis and evaluation cost are in Appx.C.
@@ 3.2 Training the Merit-Guided Prompt Optimizer
where $M_{\text{ref}}$ serves as the reference model, $M_{o}$ is the optimization model ($M_{\text{ref}}$ with adapters), $x$ denotes the input prompt in Fig.10, $c$ and $r$ are the chosen and rejected prompts defined in the DPO algorithm.
@@ 4 Experiments
We evaluate the effectiveness of MePO from the following perspectives: (1) Comparison with SOTA local prompt optimizers (Sec. 4.2); (2) Evaluation on instruction-following datasets (Sec. 4.3); (3) Ablation studies on training (Sec. 4.4) and merits (Sec. 4.5); and (4) Further analysis (Sec. 4.6)
@@ 4.1 Setups
MePO is trained using Qwen2.5-7B-Instruct2, selected for its strong generation ability among similarly sized models Qwen et al. (2025). This choice ensures that the optimizer is built on a lightweight LLM that maximizes performance within its size class, allowing efficient deployment without sacrificing generation quality.
Datasets We evaluate MePO on a diverse set of tasks, including ARC (Easy and Challenge) (Clark et al., 2018), GSM8K (Cobbe et al., 2021), BigBench-Hard (BBH, 25 datasets) (Suzgun et al., 2023), and PiQA (Bisk et al., 2020), as well as three instruction-following benchmarks: Vicuna Eval Chiang et al. (2023), BPO-Test Eval Cheng et al. (2024), and Self-Instruct Eval Wang et al. (2023). Following Lu et al. (2025), we adopt 3-shot for GSM8K, BBH, and PiQA, and zero-shot for others, with instructions for answer extraction. Prompt optimization is applied only to the test query, while others remain unchanged.
Baselines We compare MePO with two state-of-the-art local prompt optimization models: BPO (Cheng et al., 2024) and FIPO (Lu et al., 2025). We also use the inference model itself as an optimization baseline, motivated by the inherent alignment between a model and its own preferences, as leveraging an LLM’s natural capability to refine prompts for itself better reflects its aligned outputs (Xiang et al., 2025). We evaluate optimizers’ performance across seven inference model families: Qwen (Yang et al., 2024), Tulu (Ivison et al., 2023), Gemma Team et al. (2024), LLaMA (Touvron et al., 2023), Vicuna Zheng et al. (2023), Deepseek Guo et al. (2025) and GPT4o Achiam et al. (2023). Additional implementation details and dataset descriptions are provided in Appx.D.
@@ 4.2 Main Results
We present the performance of each optimizer, including our MePO, on different inference models. The results across QA and math datasets, summarized in Table 2, lead to the following conclusions:
Model Method Vicuna Eval BPO-test Eval Self-Instruct Eval
A B A Win Tie B Win A Win Tie B Win A Win Tie B Win $\Delta$ WR
GPT-4o MePO - 31.7 61.9 6.4 32.0 61.0 7.0 47.5 50.0 2.5 +31.8
MePO Inference 25.0 63.5 11.5 30.0 53.5 16.5 43.8 53.7 2.5 +22.8
MePO BPO 23.4 61.9 14.7 28.5 54.0 17.5 41.3 55.1 3.7 +19.1
MePO FIPO 24.2 62.3 13.5 22.5 62.0 15.5 30.0 65.0 5.0 +14.2
DeepSeek-R1 MePO - 7.5 86.1 6.4 37.0 58.0 5.0 31.2 52.5 16.3 +16.0
MePO Inference 27.8 66.3 5.9 16.5 71.0 12.5 20.7 55.0 24.3 +7.4
MePO BPO 19.8 73.0 7.2 18.5 71.0 10.5 32.5 60.0 7.5 +15.2
MePO FIPO 16.7 76.6 6.7 26.0 60.0 14.0 31.3 60.0 8.7 +14.9
Qwen-7B-chat MePO - 52.7 12.3 35.0 54.5 23.5 22.5 60.0 15.0 25.0 +28.2
MePO Inference 47.2 12.3 40.5 54.0 13.5 32.5 46.3 24.9 28.8 +15.2
MePO BPO 51.2 23.3 25.5 51.2 6.3 42.5 59.0 6.0 35.0 +19.5
MePO FIPO 45.2 32.3 22.5 49.0 11.8 39.2 49.9 11.4 38.7 +14.6
Vicuna-7B MePO - 56.8 10.7 32.5 59.0 9.5 31.5 61.3 8.7 30.0 +27.7
MePO Inference 48.8 18.7 32.5 56.5 12.5 31.0 46.3 23.7 30.0 +19.4
MePO BPO 42.0 18.3 39.7 51.0 12.0 37.0 48.8 22.5 28.7 +12.1
MePO FIPO 36.9 33.3 29.8 66.5 15.0 18.5 53.8 12.5 33.7 +25.1
[…]
Efficacy MePO consistently outperforms all baselines across datasets and task models, with average accuracy gains of 1% on Qwen2-7B, 2.37% on Tulu2-7B, 2.34% on LLaMA2-7B, 2.86% on Gemma2-9B and 1.69% on Qwen3-8B. Notable improvements are observed on GSM8K with LLaMA2-7B and Gemma2-9b (+4.17%, +3.31%), ARC-Easy and BBH with Tulu2-7B (+4.5%, +3.69%) and PiQA on Qwen3-8b (+3.21%).
Robustness Inference models often perform better when the optimizer shares the same architecture, reflecting alignment with model-specific preferences. For example, BPO (based on LLaMA2) performs better with LLaMA2-7B, while FIPO (based on Tulu2) excels with Tulu2-7B. However, both optimizers show misalignment when responding under different architectures such as Gemma2, Qwen2 and Qwen3. Compared with inference model itself as the optimizer, on GSM8K, the inference model outperforms BPO by +2.2% (Gemma2), +1.63% (Qwen2), and +3.11% (Qwen3), and surpasses FIPO by +6.52% on Qwen3. On ARC-Easy, it outperforms BPO by +0.33% (Qwen2) and +1.1% (Qwen3), and FIPO by +0.73% (Gemma2) and +5.32% (Qwen3). On PiQA, Qwen3 surpasses BPO by +0.76% and FIPO by +6.15%. Both BPO and FIPO fail to transfer effectively to Qwen3, with average performance drops of 0.8% and 3.97%, respectively. In contrast, MePO consistently outperforms inference model optimization across all datasets, indicating that its learned prompt merits are broadly interpretable across architectures, demonstrating both robustness and strong generalization capability.
Further results on DeepSeek models and significance tests are in Appx. E.3-E.7.
4.3 Instruction-following Dataset Evaluation
To more accurately evaluate alignment quality, we assess three instruction-following datasets across four widely used inference models: GPT-4o, DeepSeek-R1, Qwen-7B-Chat [^7] https://huggingface.co/Qwen/Qwen-7B-Chat, and Vicuna-7B [^8] https://huggingface.co/lmsys/vicuna-7b-v1.5. We compare MePO with three optimizers: the inference model itself, BPO, and FIPO. All comparisons are scored by GPT-4o to ensure consistency and high-quality preference judgments, following Liang et al. (2024).
As shown in Table 3, MePO outperforms all compared methods across all datasets. Even for powerful inference models such as GPT-4o and DeepSeek-R1, the clear, merit-aligned structure of MePO’s optimized prompts leads to further improvements. The results also suggest that stronger models can benefit from prompt optimization, as their superior generation capabilities enable them to better leverage high-quality prompts. Additional comparisons with POIR are provided in Sec. E.6.
4.4 Ablation Study: Training Ablation
Task Model Optimizer ARC-E ARC-C GSM8K BBH PiQA Avg.